OpenStudy (anonymous):

A small train at an amusement park is carrying passengers. One-fourth of the passengers are 5 years old, 1/2 of the passengers are 6 years old, 1/8 are 7 years old, and 1/8 are 9 years old. What is the expected age for one of the passengers on the train? 6 years 6.5 years 6.25 years 6.75 years

OpenStudy (misty1212):

hi dear you have something missing here

OpenStudy (misty1212):

in your question you have one fourth are 5, but not how many are 6, 7 or 9

OpenStudy (anonymous):

im sorry ill fix it lol thankss

OpenStudy (anonymous):

What did you choose as your answer?

OpenStudy (misty1212):

what ever it is, multiply and add \[\frac{1}{4}\times 5+whatever \times 6+whatever \times 7+whatever\times 9\]

OpenStudy (misty1212):

\[\frac{1}{4}\times 5+\frac{1}{2} \times 6+\frac{1}{8} \times 7+\frac{1}{8}\times 9\]

OpenStudy (anonymous):

thanks so much got answer :)

OpenStudy (misty1212):

glad you got it